### v2

# Dataset Awareness is not Enough: Implementing Sample-level Tail Encouragement in Long-tailed Self-supervised Learning

Abstract
Self-supervised learning (SSL) has shown remarkable data representation capabilities across a wide range of datasets. However, when applied to real-world datasets with long-tailed distributions, performance on multiple downstream tasks degrades significantly. Recently, the community has begun to focus more on self-supervised long-tailed learning. Some works attempt to transfer temperature mechanisms to self-supervised learning or use category-space uniformity constraints to balance the representation of different categories in the embedding space to fight against long-tail distributions. However, most of these approaches focus on the joint optimization of all samples in the dataset or on constraining the category distribution, with little attention given to whether each individual sample is optimally guided during training. To address this issue, we propose Temperature Auxiliary Sample-level Encouragement (TASE). We introduce pseudo-labels into self-supervised long-tailed learning, utilizing pseudo-label information to drive a dynamic temperature and re-weighting strategy. Specifically, We assign an optimal temperature parameter to each sample. Additionally, we analyze the lack of quantity awareness in the temperature parameter and use re-weighting to compensate for this deficiency, thereby achieving optimal training patterns at the sample level. Comprehensive experimental results on six benchmarks across three datasets demonstrate that our method achieves outstanding performance in improving long-tail recognition, while also exhibiting high robustness.
1 Introduction
Figure 1: Comparison between SimCLR [5] baseline trained on balanced CIFAR10[25] and imbalanced CIFAR10-LT [11] and our TASE trained on CIFAR10-LT. Three Splits (Head, Mid and Tail) are shown as different colours (blue, red and green) respectively [29]. We also report the average accuracy and range of variation over ten classes.
Self-supervised learning (SSL) [17, 18], which aims to learn robust and generic feature representations from unlabeled data, has been consistently explored in recent years [7, 3, 45]. Contrastive learning, a leading SSL paradigm, seeks to learn discriminative representations by encouraging the embedding similarity of augmented view pairs, and has achieved remarkable success across several domains [46, 15]. Nevertheless, most contrastive learning methods are conducted on curated datasets like ImageNet [12]. In contrast, real-world data distributions are mainly characterized by long-tailed distributions [34], which could cause a significant degradation in performance, as illustrated in Fig. 1. On the relatively simple CIFAR10-LT dataset, the baseline model demonstrates a notable performance collapse. For certain tail classes, the accuracy even falls below 30%, indicating that SSL still suffers from the long-tailed problem. The difficulty of detecting or modeling such implicit imbalance is a significant obstacle to migrating SSL to large-scale, uncurated, real-world datasets.
In recent years, for the supervised paradigm, the long-tailed problem has been extensively studied from multiple perspectives, deriving strategies such as re-weighting [53], re-sampling, and ensemble learning, which have achieved significant success. In terms of re-weighting, a common approach is to increase the weight proportion of tail-class samples, in order to occupy a larger gradient and encourage the model to focus on tail-class samples. Turning to self-supervised learning, particularly contrastive learning, some works have proposed using temperature assignment to counteract the effects of long-tailed distributions [41, 26]. Specifically, models can be made to focus on different semantics by adjusting the temperature parameters of the contrastive objective. A lower temperature encourages sample features to push similar samples away as much as possible, independently forming semantic clusters. This is a pattern emphasizes category discrimination, while a higher temperature allows greater tolerance for similar samples, focusing more on semantic difference learning and instance discrimination. Such learning pattern difference is a potential way to combat long-tailed distributions.
Then our perspective shifts back to self-supervised long-tailed learning, some research has sought to adapt the aforementioned mechanisms to this domain. TS [26] attempts to employ a global temperature with cosine variation to modify the model’s learning pattern between instance discrimination and category discrimination. GH [52] emphasizes the necessity for spatial homogeneity in imbalanced learning scenarios and proposes a plug-and-play objective for harmonizing the embedding space of each category. Nevertheless, none of these efforts achieve optimal training for all samples, and intuitively, it is more efficient to ensure that each sample is appropriately trained than to constrain the global.
Motivated by the above considerations, we notice pseudo-labels assignment in semi-supervised learning, and then use pseudo-labels of each sample to achieve sample-level encouragement. Nevertheless, due to the long-tailed distribution, the tail classes with few samples are more likely to yield severe misassignments. In order to avoid serious misleading of the tail classes, several pseudo-labels assignment methods from semi-supervised learning, which use pseudo-labels as targets for consistency regularization [36, 47], cannot be directly transferred to the long-tailed scenario. Therefore, we make the following attempts and propose TASE. Specifically, we assign appropriate temperature based on the category to which the pseudo-labels of samples belong. This allows the head class to learn semantic differences from other similar samples, and while the tail class is separated from other sample clusters, thereby generating individual categories. Nevertheless, during the training process, the head and tail classes are assigned the maximum and minimum temperatures respectively, but the range of temperature parameters is often treated as a fixed parameter, which results in the strength of tail encouragement having the same intensity for datasets with different imbalance ratios. We call this as quantity awareness flaw of temperature strategy. To address this flaw, we propose re-weighting strategy to provide explicit quantity awareness. In particular, we add a weight to each sample that is inversely proportional to the number of samples in the category to which it belongs.
Compared to existing work, our method has the following advantages: 1) Compared to previous work on TS [26] and GH [52], we are able to modify the training process at the sample-level, providing the most appropriate encouragement to each sample to counteract the long-tailed distribution. 2) Compared to the approach in semi-supervised learning, where pseudo-labels are directly used as targets for consistency regularization, our method does not require pseudo-labels for direct training, and it is better at resisting the misleading effects of incorrect label assignments.
We summary contributions of this paper as below:
  • We note the drawback of existing work on long-tail self-supervised learning that it is difficult to perform tail encouragement for the sample level, and propose to introduce the idea of pseudo assignment into long-tail self-supervised learning to perform sample-specific learning for each sample.
  • We propose TASE as a preliminary attempt to implement sample-level control in long-tail self-supervised learning, assigning appropriate temperatures driven by pseudo assignments for different samples during training to achieve optimal training patterns.
  • We implement comprehensive experiment across six benchmarks on three commonly used datasets, and the experimental results demonstrate our model achieves superior performance.
2 Related Work
Supervised Long-tailed Recognition. Classical supervised long-tail recognition can be categorized into three directions: category-level re-balancing, information augmentation and module improvement [49]. Related to our work, exploration of category-level re-balancing mainly focuses on re-sampling, re-weighting and logit adjustment [21, 53, 24, 42]. Decoupling [21] explores the impact of different sampling strategies, and finds that square-root sampling [31] and progressively-balanced sampling are better suited for long-tailed datasets. LOCE [13] proposes a memory-agumented feature sampling to dynamically monitor the prediction score of each classes to drive re-sampling rate distribution. Oriented towards the supervised contrastive learning paradigm, BCL [53] develops a novel loss for balanced training, which considers the averaging of negative class gradients and the complement of mini-batch tail classes. Logit adjustment [32] utilises training label frequencies to adjust the prediction logit. VS loss [24] conducts a comprehensive analysis on the impact of additive and multiplicative logit-adjusted losses, and further formulates a novel strategy that combines the best of both worlds.
Explorations on information augmentation seek to introduce additional information into training process, mainly containing transfer learning and data augmentation [44, 50, 16, 27, 33, 14]. SSP [44] systematically analyzes the benefits of combining semi-supervised and self-supervised manners to standard learning for balanced feature space learning. SSD [27] develops a soft labels generator driven by self-supervised distillation, producing robust auxiliary information to transfer knowledge from long-tailed distribution. DeiT-LT [33] aims to tackle the problems of high efficiency training from scratch, introducing an effective distillation construct from a flat CNN teacher via CLS token and distillation DIST token to support a head expert and a tail expert, respectively. Remix [8] benefits from re-balanced data mixup to boost long-tail learning. SBCL [19] employs subclass-balancing adaptive clustering and bi-granularity contrastive loss to facilitate the implicit mining of sub-classes and optimise the balance of sub-classes.
Module improvement can be divided into classifier design and ensemble learning [10, 50, 22, 43, 48]. Paco [10] introduces a set of parametric class-wise learnable centers to adaptively enhance the optimisation of latent space. RIDE [43] leverages a strategy of multiple experts driven by KL-divergence based loss to benefit hard example learning. SADE [48] explores a multi-expert scheme to handle test-agnostic long-tailed recognition, innovating diversity-promoting expertise-guided losses to adaptively aggregate experts for handling unknown distribution.
Self-supervised Long-tailed Recognition expects to automatically correct potential category bias without ground-truth labels and data distributions. There are several recent explorations for this purpose. SDCLR [20] constructs a pruned dynamic self-competitor for mining forgettable samples, which are then implicitly emphasized in the contrastive process. DnC [38] alternates between base contrastive learning and multi-experts hard negative mining, aggregating features from a distillation step to mitigate the collapse of the rare classes. TS [26] leverages an oscillating temperature factor with a cosine schedule to dynamically learn both group-wise and instance-specific features. FASSL [28] presents a frequency-aware prototype learning to identify prototypes guidance reflecting latent distribution from unlabeled data, benefiting re-balanced learning. From a data perspective, BCL-I [51] exploits the memorization effect of DNNs to automatically drive an augmentation of each instance, thereby aiding the learning of discriminative features. GH [52] develops a novel contrasting loss from the perspective of geometric harmonisation, dynamically approaching the class-level uniformity of the latent space.
Pseudo-labels Assignment. DeepCluster [2] leverages the cluster assignments as pseudo-labels, driving an end-to-end unsupervised learning with the soft labels classification pretext task. Subsequently, a large number of methods have been developed to address degenerate solutions, i.e., to prevent collapse. Unlike end-to-end approaches, SCAN [40] employs a method that decouples representation learning from clustering, ensuring that the clustering process does not depend on low-level features. SeLa [1], on the other hand, focuses on maximizing the information between labels and input data indices. SwAV [4] develops a method that maps images to a set of trainable prototype vectors, which simultaneously avoids the computational cost of pairwise comparisons and improves the model’s scalability and training efficiency.
3 Method
3.1 Preliminary
For a given dataset $\mathcal{D}$, every instance is denoted by $(\bm{x},\bm{y})\in\mathcal{D}$,where $\bm{y}$ is corresponding class label of input $\bm{x}$. Contrastive learning expects to close similarity of positive pairs $(\bm{x_{i}},\bm{x_{i^{\prime}}})$ produced by a random augmentation and repel similarity of negative pairs $(\bm{x_{i}},\bm{x_{j}})$. The contrastive loss for instance $i$ is defined by:
$\ell_{i}=-\log\frac{S(\bm{v}_{i},\bm{v}_{i^{\prime}},\tau)}{S(\bm{v}_{i},\bm{v}_{i^{\prime}},\tau)+\sum_{v_{j}\in\mathcal{B}^{-}_{i}}S(\bm{v}_{i},\bm{v}_{j},\tau)}~{}$ (1)
$S(\bm{v}_{i},\bm{v}_{j},\tau)=\exp(sim(\bm{v}_{i},\bm{v}_{j})/\tau)~{}$ (2)
Where $sim(\cdot)$ denotes the cosine similarity, $B^{-}_{i}$ denotes a set of negative samples of $x_{i}$ and $v_{i}=g(f(x_{i}))$ is embedding of $x_{i}$. $f(\cdot)$ and $g(\cdot)$ are encoder and projection head network, respectively.
3.2 Temperature Property
In recent years, research has been conducted to examine the mechanisms of different temperature parameters in greater details [35, 41]. Several evidence from both quantitative and qualitative experiments has demonstrated that temperature parameters exert a profound influence on the uniformity and tolerance of training process. Previous work [41] calls these two opposing properties as the uniformity-tolerance dilemma, corresponding to instance-discrimination and category-discrimination, respectively. To be specific, for Eq. 1 and Eq. 2, given a small $\tau$, the penalty will be dominated by hard negative samples, and therefore more attention will be paid to push the more similar samples farther away, i.e., the tolerance of this feature point is small. Accordingly, when the tolerance of all the points is small, the latent space will be paved over as much as possible, presenting a high uniformity feature. In contrast, a large $\tau$ exhibits high tolerance and low uniformity; therefore, feature points will try to occupy as little space as possible.
We give a gradient analysis to illustrate this. The gradient for the similarity between $v_{i}$ and a random negative sample $v_{k}$ is denoted as Eq. 3, and the ratio of its gradient to the sum of gradients with respect to all negative samples is given by Eq. 4. When $sim(\bm{v}_{i},\bm{v}_{k})$ tends to 1, i.e., $v_{k}$ is a hard negative example, with the value of $\tau$ linearly decrease, $r_{k}$ will increase significantly, implying that the hard negative samples gradually dominates gradient direction.
$\frac{\partial\mathcal{L}(x_{i})}{\partial sim(\bm{v}_{i},\bm{v}_{k})}=\frac{1}{\tau}\frac{S(\bm{v}_{i},\bm{v}_{k},\tau)}{S(\bm{v}_{i},\bm{v}_{i^{\prime}},\tau)+\sum_{\bm{v}_{j}\in\mathcal{B}_{i}^{-}}S(\bm{v}_{i},\bm{v}_{j},\tau)}$ (3)
$r_{k}=\frac{\exp(sim(\bm{v}_{i},\bm{v}_{k}))/\tau)}{\sum_{\bm{v}_{j}\in\mathcal{B}_{i}^{-}}\exp(sim(\bm{v}_{i},\bm{v}_{j})/\tau)}$ (4)
3.3 Structural Separability and Category Uniformity
In this section, we provide an analysis about the expectation of the semantic spatial structure and how the dynamic temperature strategy could help reach this ideal structure.
In order to obtain robust downstream task performance, higher requirements are demanded for the semantic space structure of the data representation, which we describe as structural separability and category uniformity. As shown in Fig. 2, structural separability expects that samples from different categories form category semantic clusters and maintain strong linear separability. At the same time, in order to focus on the tail categories, category uniformity expects that each category occupying embedding space should be similar, irrespective of the number.
From the perspective of temperature setting, a lower temperature can make the sample pay more attention to strong negative examples and pursue better separation. On the other hand, stronger uniformity can make the latent space occupied by each category larger, which is more suitable for categories with fewer samples. Conversely, we hope to set a higher temperature for categories with more samples, encouraging the model to learn the inherent semantic differences between samples and maintain high generalization across a variety of transfer learning tasks. To this end, we utilize a temporary feature clustering module to assign dynamic pseudo-labels as auxiliary information to guide the assignment of temperature. If the sample belongs to the head class, a higher temperature is set, and if it belongs to the tail class, a lower temperature is assigned. There is no restriction on the clustering method. We use a simple K-means to verify our idea.
Figure 2: Illustration of the structure of the desired embedding space. We use different icons in different colors to represent the samples from different categories and indicate the size of the space occupied by each category. We expect the clusters to be at a certain distance from each other and to occupy a similar amount of space.
3.4 Explicit Quantity Awareness
In Sec. 3.3, we describe how we use dynamic temperature tuning strategy to optimize spatial structure. However, we find this to be a flaw in practice, as the general temperature settings are treated as fixed hyperparameters, and pseudo-labels cannot provide information on imbalance ratios. In other words, this flaw gives rise to the fact that the same temperatures will always be assigned to the head class samples on different datasets with different imbalanced ratio. Therefore, we require an adjunct that can take advantage of the information on the distribution of pseudo-labels to jointly serve as a hybrid objective.
Therefore, we introduce a classic supervised re-weighting strategy to assist category uniformization. For each sample, as negative examples, a weight is added inversely proportional to the number of samples in the category to which they belong. And to prevent overcorrection we use the inverse of the square root of the sample size as the weights.
3.5 Training via TASE
In this section, we introduce our design of loss function feature and clustering component.
$\ell_{i}=-\log\frac{S(\bm{v}_{i},\bm{v}_{i^{\prime}},\tau_{i})}{S(\bm{v}_{i},\bm{v}_{i^{\prime}},\tau_{i})+\sum_{\bm{v}_{j}\in\mathcal{B}^{-}_{i}}\frac{1}{\sqrt{|D_{j}|}}S(\bm{v}_{i},\bm{v}_{j},\tau_{i})}~{}$ (5)
$\mathcal{L}_{TASE}=\frac{1}{|D|}\sum_{\bm{v}_{i}\in D}{\ell_{i}}$ (6)
First, we give our hybrid cross-paradigm loss as Eq. 6, where $\tau_{i}$ denotes a dynamic temperature assigned by temporary labels, and $|D_{j}|$ denotes the number of the samples that is allocated to the same class as $v_{j}$.
Our training is divided into two stages: warming for $B$ epochs and cluster balance training for the rest epochs. In the warming phase, we utilize original SimCLR to seek a good cluster initialization. After $B$ epochs, we use a progressive temperature factor until $S$ epochs.
The pseudo-code is summarized as Algorithm 1
Algorithm 1 Pseudo-code of our proposed TASE.
Input: dataset $\mathcal{D}$, number of epochs $E$, number of warming epochs $B$, cluster frequency $F$, encoder network $f$
Output: pretrained model parameter $\theta_{E}$
Initialize: model parameter $\theta_{0}$
1:  Warm up model $\theta_{0}$ for $B$ epochs via SimCLR.
2:  Initialize the pseudo-label allocation $M$.
3:  for epoch $e=B,B+1,\ldots,E$ do
4:     Compute the distribution of temporary classes assignment $P$ by $M$.
5:     for mini-batch $n=1,2,\ldots,N$ do
6:        Obtain the temperature $\mathbf{t}$ for each sample by $P$
7:        Compute the proposed $\mathcal{L}_{\mathrm{TASE}}$.
8:        Uptate model $\theta$ by $\mathcal{L}_{\mathrm{TASE}}$.
9:     end for
10:     if epoch $e$ $F$ == 0 then
11:        Update pseudo-label allocation $M$.
12:     end if
13:  end for
3.6 More Discussion
Computational Complexity and Storage Consumption. Our method maintains a temporary list of pseudo-labels with length equal to the number of samples and performs a cross-batch clustering operation once in a certain number of rounds. Our method requires only a small increase in storage consumption and computational overheads.
Expandability Our TASE can be combined with other contrastive learning frameworks [15, 17] or non-contrastive self-supervised frameworks [18]. Correcting potential imbalances by using only temporary image representations results in high availability and low overhead. The feasibility and extensibility of TASE have been validated by combining it with SimCLR, and we will make combining it with other self-supervised learning frameworks in the future.
4 Experiment
4.1 Datasets
CIFAR10-LT and CIFAR100-LT [11] are long-tailed versions of the CIFAR10 and CIFAR100 datasets, respectively. The original CIFAR10 [25] dataset comprises of 32×32 images from 10 classes, with 50k images allocated for training and 10k images for testing. The imbalance factor is defined as the number of the head classes divided by the number of tail classes. Following SDCLR [20] and TS [26], we set the imbalance factor to 100. Each class contains between 4,500 and 45 images according to the Pareto distribution. In total, there are 11,165 images in the training and test sets, with 10,000 images remaining.
Similarly, the CIFAR100 dataset has 100 classes of images with 50k images for training and 10k for testing. The number of images per class for training in CIFAR100-LT varies from 450 to 4, with a total of 9,754 images. We average all of our results over three runs across five splits to reduce randomness to the greatest extent possible.
ImageNet100-LT is a subset of ImageNet100 [37] dataset, consisting of 12.21k images from 100 classes for training and 5k images for testing. The number of images of each class in the training set range from 1280 to 5 according to the Pareto distribution.
4.2 Experimental Setup
Evaluation. TS [26] provides a comprehensive validation that employs K nearest neighbours [9] (KNN) to verify the latent space structure by directly evaluating the learned local representations, and two linear probing (LP) methods to validate the generalization and linear separability [20] of learned embedding, which we call the balanced minimum sampling linear probing (MS LP) and the long-tailed linear probing (LT LP). We follow their validation methods, and in addition, we provide two other more widely used validation methods, the 1% sampling linear probing (1%S LP) and the full set linear probing (Full LP), all of which we will explain in detail below.
KNN [9] can directly evaluate the learned spatial structure. We compute L2-normalized distances between training samples and testing samples. The class of each testing image is then predicted as the majority class among the top-K closest training samples. Two series of accuracy have been established for KNN with $K=1$ and $K=10$ marked as KNN@1 and KNN@10, respectively.
Linear probing [23] trains a linear classifier layer on the top of pre-trained feature extractor by using specified data and then evaluates accuracy on the testing set. By default, we use full set as fine-tune dataset and validate it on the testing set called Full LP. Considering practical scenarios, we also consider two balanced few shot setups [6]: minimum sampling linear probing (MS LP) and 1% sample linear probing (1%S LP), both use 1% of the full dataset and have an equal number of samples between each class. The difference is that the former samples the data in the pre-train dataset as a fine-tuning dataset with better uniformity, while the latter randomly samples 1% of all the dataset, which is more closely aligned with the actual scenario. Additionally, we also conduct evaluation driven by imbalanced fine-tuning dataset called long-tailed linear probing (LT LP), using original imbalanced training set.
Method KNN@1 KNN@10 MS LP 1%S LP LT LP Full LP
ResNet18
SimCLR 59.84 60.19 68.29 62.98 61.86 75.37
SDCLR 61.92 60.63 71.74 71.74 65.38 80.49
TS 63.09 62.91 71.86 71.29 65.03 79.08
FASSL 68.54 80.69
TASE 64.06 63.70 72.78 72.60 65.46 80.35
ResNet50
SimCLR 62.51 61.49 72.83 72.17 65.78 83.75
SDCLR 62.59 61.27 73.30 72.73 67.07 85.59
TS 63.11 61.94 73.54 73.28 66.80 84.76
TASE 64.17 62.83 74.16 74.27 68.15 85.46
Table 1: Performance comparison on CIFAR10-LT dataset. We utilizes Resnet18 and Resnet50 as backbone to reproduce some recent long-tailed self-supervised methods, and we bold the best performance.
Implementation Details. Following previous works [20, 26], in the case of smaller-scale datasets CIFAR10-LT and CIFAR100-LT, the ResNet-18 architecture is employed as the backbone for training over 2000 epochs, while for the ImageNet100-LT dataset, the ResNet-50 is utilised for training over 800 epochs. Additionally, we train our model with SGD optimizer, batchsize of 512 for CIFAR10-LT and CIFAR100-LT, 256 for ImageNet100-LT, momentum of 0.9, feature dimension of 128 and weight decay of $1e^{-4}$. We adopt a learning rate followed by cosine annealing schedule with 10 epochs linear warm-up and set learning rate as 0.5 for CIFAR and 2 for ImageNet.
As for our unique hyperparameters, The value of clusters $K$ is set as equal to the number of classes contained in the dataset, i.e., $K$=10 for CIFAR10-LT and 100 for ImageNet100-LT and CIFAR100-LT. The value of B is equivalent to the starting epoch of clustering, i.e. B=50 for CIFAR10-LT and ImageNet100-LT, 300 for CIFAR100-LT. In all three datasets, we assign the frequent of clustering F to 10. Epochs of warming S=500 for CIFAR10-LT, 1000 for CIFAR100-LT, and 200 for ImageNet100-LT. We will have an in-depth analysis of parameters sensitivity and detailed experimental results in Sec. 4.4.
To gain a more detailed insight of the quantitative impact of our method for calibrating domination of majority classes, we divide all classes of each dataset into three groups from highest to lowest sample size: Head, Mid and Tail. For CIFAR10-LT and CIFAR100-LT datasets, we split the three groups evenly, 4:3:3 for CIFAR10-LT and 34:33:33 for CIFAR100-LT. As for ImageNet100-LT dataset, following [29], defining head classes as classes with over 100 samples, and tail as classes under 20 samples, classes have the number of samples from 20 to 100 are assigned as Mid classes.
4.3 Experimental Results
In this section, we summarize the comprehensive experimental results by comparing our TASE with recent methods across a range of benchmarks, and offer insights into the observed outcomes.
In the CIFAR10-LT and CIFAR100-LT datasets, we conduct comparison experiments across all six benchmarks previously mentioned. As shown in Tab. 1 and Tab. 2, for comparison of latent space structure with the ResNet18 backbone model, i.e., KNN@1 and KNN@10, we achieve a significant consistent improvement compared to previous methods, which demonstrates that our approach efficiently optimises local embedding representations, which are based on the learned metric. In terms of practical considerations, we adopt three settings of linear probing experiment: 1%S LP, MS LP and LT LP. We find that the average accuracy is positively correlated with the number of fine-tuned samples. For the MS LP and 1%S LP benchmarks, the proposed method continues to demonstrate satisfactory performance. Differences in data intersections for training and fine-tuning yielded by the two settings exerts the most significant influence on SimCLR, while has a limit impact on other methods. At setting of LT LP, we utilize original long-tailed pre-training set for fine-tune. Compared to MS LP and 1%S LP, LT LP has a greater uniformity and more fine-tuning instances. Despite the reduction in the lead, our TASE still achieves a competitive result. But for the settings of Full LP, our TASE does not perform as expected, especially in the CIFAR100-LT dataset.
Method KNN@1 KNN@10 MS LP 1%S LP LT LP Full LP
ResNet18
SimCLR 28.81 28.12 25.70 22.51 31.2 49.76
SDCLR 30.19 29.59 26.63 27.32 34.38 55.00
FASSL 27.11 55.27
TS 31.06 30.06 28.89 28.32 33.28 51.64
GH 29.34 28.58 27.73 27.60 34.13 53.18
BCL-I 30.52 30.07 28.62 28.49 34.99 55.79
TASE 31.27 30.81 29.02 28.69 33.96 52.21
ResNet50
SDCLR 30.17 29.57 26.15 25.71 35.67 59.43
TS 31.50 30.60 28.46 28.74 36.00 58.81
BCL-I 30.38 29.68 28.68 29.07 36.62 59.30
TASE 32.14 31.34 29.19 29.17 36.73 59.57
Table 2: Performance comparison on CIFAR100-LT dataset. We utilizes Resnet18 and Resnet50 as backbone to reproduce some recent long-tailed self-supervised methods, and we bold the best performance.
Method K@1 K@10 1%S LP LT LP Full LP
SimCLR 37.10 38.00 42.16 44.82 65.46
SDCLR 36.06 37.36 42.38 46.40 66.48
TS 38.36 38.86 45.18 47.26 67.82
TASE 39.54 40.26 45.48 47.82 68.30
Table 3: Performance comparison on ImageNet100-LT dataset. The best performance is highlighted in bold.
In order to gain a deeper understanding of whether this performance discrepancy persists under different alternative models, a series of analogous experiments were conducted around ResNet50. We summarise the experimental results in the ResNet50 section of Tab. 1 and Tab. 2. The experimental results indicate that our method does not result in performance collapse when utilising ResNet50, and maintains an excellent performance like KNN series in settings of ResNet18 backbone. It is worth noting that both temperature adjustment methods (i.e., TS and TASE) yield the similar performance collapse, and this crash appears to be exclusive to the CIFAR100-LT dataset, which features a greater number of categories, and was more pronounced when utilising more fine-tuned data.
$\bm{K}$ K@1 K@10 MS LP 1%S LP LT LP Full LP
70 30.46 29.46 27.37 29.41 33.23 52.54
100 30.76 29.31 28.54 28.45 32.80 51.82
130 30.66 29.19 27.47 28.51 33.53 52.40
200 30.69 29.36 27.31 28.46 33.27 51.84
$R$ 0.3 0.27 1.23 0.96 0.73 0.72
Table 4: Effect of $\bm{K}$ on CIFAR100-LT dataset with ResNet18. $\bm{K}$ denotes the number of clusters. The best performance is highlighted in bold, while the worst performance is underlined. Additionally, the range of fluctuations has been indicated with $R$.
Further, we also summary performance comparison on ImageNet100-LT dataset with ResNet50, as shown in Tab. 3.
$\bm{B}$ K@1 K@10 MS LP 1%S LP LT LP Full LP
0 62.77 61.50 71.61 71.29 63.58 79.72
50 63.08 62.30 71.78 70.84 64.50 79.58
100 63.01 61.67 71.10 70.55 63.16 79.29
300 63.36 62.04 71.40 70.19 63.46 79.76
500 63.14 61.82 71.36 70.77 64.43 79.35
$R$ 0.59 0.8 0.68 1.1 1.34 0.47
Table 5: Effect of $\bm{B}$ on CIFAR10-LT dataset with ResNet18. $\bm{B}$ denotes a hyperparameter that determines when clustering begins. The best performance is highlighted in bold, while the worst performance is underlined. Additionally, the range of fluctuations has been indicated with $R$.
4.4 Analysis of Sensitivity
We verify the hyperparameters sensitivity of our TASE. Firstly, as we simply utilize K-means [30] for assigning the dynamic labels, the number of clusters $\bm{K}$ should be given carefully consideration. It is challenging to ascertain the total number of image categories without ground truth labels, and finding the optimal parameters is computationally expensive and time-consuming working on large-scale datasets. However, the strong robustness exhibited by proposed TASE will alleviate this problem, as shown in Tab. 4. We validate performance of various $\bm{K}$ on CIFAR100-LT dataset, the best item is indicated by bold text and underline for the worst one. Additionally, we report the accuracy range of each benchmark, spanning from 0.27% to 1.23%. Our method maintains competitive results even with K=200, and the high robustness of $K$ helps us to find the optimal parameters more efficiently in the face of unknown datasets.
Furthermore, we further conduct sensitivity experiment of $\bm{B}$, which determines when to start using the clustering results to influence the learning objective. Setting $B$ larger may harm the concentration on long-tailed learning process, while too small may have impact on the clustering results and drive the model to learn from non-ideal metrics, as shown in Tab. 5. We fix $K$=10 for experiments on the CIFAR10-LT dataset (split 1). From the results, we can see that the performance difference in $B$ variation from 0 to 500 is minimal with the setup of 2000 epochs training.
Method $\bm{\tau}$ $\bm{w}$ KNN@1 KNN@10 1%S LP MS LP LT LP Full LP
[…]
Table 6: Ablation Experiments. Performance comparison on CIFAR10-LT dataset (Split 1) with ResNet18. $\bm{\tau}$ denotes employing a dynamic temperature strategy, and $\bm{w}$ denotes using a re-weight operation. The best performance is highlighted in bold.
$\bm{F}$ $\bm{S}$ 200 500 800 1000 $R_{F}$
1 71.49 70.39 71.16 71.13 1.1
5 70.22 70.29 70.14 70.23 0.15
10 70.70 70.84 70.65 70.48 0.36
30 70.97 71.21 70.54 70.97 0.67
$R_{S}$ 1.27 0.92 1.02 0.90
Table 7: The 1%S LP performance of varying $\bm{F}$ and $\bm{S}$ on CIFAR10-LT dataset with ResNet18. $\bm{F}$ denotes the frequent of clustering, while $\bm{S}$ denotes epochs of warming. The best performance is highlighted in bold. $R_{F}$ and $R_{S}$ denote the range of fluctuations of accuracy with fixed $\bm{F}$ and $\bm{S}$, respectively.
Beyond that, we also conduct more sensitivity validation of clustering hyper-parameters shown as Tab. 7, i.e., clustering frequency $F$ and clustering warming epochs $S$. We report the 1%S LP performance on CIFAR10-LT split 1 with $K$=10 and $B$=50, and bold the best accuracy. The results demonstrate that the performance remained consistent across a range of parameter combinations. The proposed TASE displays a great robustness across multiple settings, thereby reducing the additional tuning overhead typically associated with transferring to other public datasets.
4.5 Ablation Study
To verify the importance of each component of proposed TASE, we conduct ablation experiments for proposed component, as shown in Tab. 6. Our approach has adaptive temperature awareness and quantity-aware re-weighting, both of which utilise the temporary labels obtained from clustering to correct the neglect of tail classes.
Compared to SimCLR, our method maintains a consist gain from clustering information. TASE improves the performances on all the benchmarks ranging from 3.97% to 9.26%, with an average improvement of 6.75%. Even only utilising the re-weighting strategy, TASE can also have an average of 3.90% improvement.
In contrast to the TS using temperature cosine variation strategy, TASE still improves accuracy by about 0.4% on each benchmark when only using temperature adaptation. These results corroborate the effectiveness of our proposed clustering enhancement approach.
SDCLR [20]
TASE (Ours)
Figure 3: T-SNE Visualization. Different colors indicate different classes.
4.6 T-SNE Visualization
T-SNE provides a qualitative perspective into the learned discriminative representation. As shown in Fig. 3, we select samples from all 10 classes of CIFAR10-LT to generate plots. Compared to the SDCLR, our TASE exhibits tighter clusters and more clearly distinguishing boundaries. The results demonstrate that our method is effective in producing high-quality auxiliary supervision when facing long-tailed distributions.
5 Conclusion
In this paper, we proposed a novel method for self-supervised long-tailed recognition called TASE. We introduce pseudo-label assignment to drive the dynamic temperature strategy and re-weighting to counteract the long-tailed distribution, training different samples with optimal training modes to obtain robust representations. Our approach demonstrates that the experience of supervised long-tail learning can be migrated to self-supervised long-tail learning with low overhead, in contrast to previous work that proposed non-generic solutions due to the absence of labels. We conduct various benchmarks on multiple datasets to demonstrate the efficiency and robustness of our proposed method. T-SNE [39] visualization analysis also provides a perspective showing better separability of our local representations.
Supplementary Material
6 Other Experimental Results
6.1 Fine-grained Results
We also summarize a detailed results of CIFAR10-LT and ImageNet100-LT to verify the performance of fine-grained classes in Tab. 10 and Tab. 11.
We provide accuracy for different categories.The situation of split is referenced in Sec. 4.2.
6.2 Results on Balanced Datasets
We also performed the same experiments on the balanced dataset, as shown in Tab. 8.
The results show that our method can achieve a significant lead on most benchmarks, and a similar performance to the previous work on Full LP. This indicates that our method is effective in identifying potential subclasses within a balanced dataset, thereby enhancing the performance.
Method K@1 K@10 MS LP 1%S LP LT LP Full LP
SimCLR 53.14 47.15 67.86 74.81 61.85 80.69
SDCLR 64.01 60.50 76.38 75.81 71.42 84.75
TS 65.33 62.95 76.99 78.56 69.70 83.32
TASE 69.61 67.44 78.27 79.02 71.77 84.15
Table 8: Performance comparison on balanced CIFAR10 dataset with ResNet18.
6.3 Results via Other Clustering Methods
Alternative clustering methods are employed to implement our TASE, result is summarized in Tab. 9. Experimental results indicate that K-means and agglomerative clustering exhibit comparable performance. More clustering methods such as discriminative clustering will be explored in future work.
Method K@1 K@10 MS LP 1%S LP LT LP Full LP
DBSCAN 60.96 60.77 69.74 68.97 61.83 78.67
A. C. 62.22 61.49 71.57 71.60 64.39 79.46
K-means 63.08 62.30 71.78 70.84 64.50 79.58
Table 9: Effect of different clustering methods on CIFAR10-LT dataset (Split 1) with ResNet18. A. C. denotes Agglomerative Clustering.
7 Datasets Source
We used three datasets: CIFAR10-LT, CIFAR100-LT, and ImageNet100-LT. They were all obtained by long-tail subset sampling from the original datasets, namely CIFAR10, CIFAR100, and ImageNet. Sources of original datasets can be found in Tab. 12. These datasets are publicly available and widely used. All data are in full compliance with all applicable ethical standards.
8 Computing Infrastructure
Our codes are based on Pytorch 2.4.0. Our models and reproductions of previous work are trained on GeForce RTX 3090 and GeForce RTX 4090.
KNN@1 KNN@10
Method Head Mid Tail All Head Mid Tail All
SimCLR 87.68$\pm$4.28 62.46$\pm$12.35 29.01$\pm$15.3 62.51 90.68$\pm$4.09 62.14$\pm$12.77 21.91$\pm$18.21 61.49
SDCLR 86.92$\pm$4.25 62.36$\pm$12.15 30.38$\pm$13.37 62.59 90.71$\pm$4.24 61.50$\pm$12.49 21.78$\pm$16.83 61.27
TS 88.39$\pm$3.99 61.91$\pm$13.02 30.59$\pm$14.46 63.11 91.46$\pm$3.84 61.26$\pm$12.60 23.26$\pm$17.07 61.94
TASE 88.63$\pm$3.85 64.15$\pm$11.74 31.58$\pm$14.09 64.17 91.86$\pm$3.66 63.11$\pm$12.03 23.84$\pm$17.50 62.83
MS LP 1%S LP
Method Head Mid Tail All Head Mid Tail All
SimCLR 79.48$\pm$5.30 70.49$\pm$11.73 66.32$\pm$12.48 72.83 79.63$\pm$4.12 71.31$\pm$14.46 63.09$\pm$11.5 72.17
SDCLR 77.75$\pm$6.21 73.10$\pm$10.82 67.57$\pm$13.71 73.30 77.11$\pm$5.97 71.91$\pm$13.98 67.71$\pm$11.34 72.73
TS 79.15$\pm$9.23 73.65$\pm$12.08 65.97$\pm$15.20 73.54 79.38$\pm$5.72 72.51$\pm$13.32 65.93$\pm$11.21 73.28
TASE 79.61$\pm$7.34 73.27$\pm$11.52 67.79$\pm$10.43 74.16 80.30$\pm$5.23 73.53$\pm$12.57 66.96$\pm$12.41 74.27
LT LP Full LP
Method Head Mid Tail All Head Mid Tail All
SimCLR 93.04$\pm$2.95 66.89$\pm$11.95 28.34$\pm$14.20 65.78 87.60$\pm$4.67 83.35$\pm$6.92 79.01$\pm$7.77 83.75
SDCLR 93.68$\pm$3.24 69.53$\pm$9.80 29.15$\pm$13.37 67.07 89.50$\pm$3.42 85.20$\pm$6.00 80.77$\pm$7.01 85.59
TS 93.70$\pm$2.73 68.98$\pm$11.82 28.75$\pm$12.30 66.80 88.00$\pm$4.54 84.41$\pm$6.65 80.79$\pm$6.31 84.76
TASE 94.14$\pm$2.52 70.77$\pm$9.30 30.86$\pm$12.94 68.15 89.03$\pm$4.23 85.03$\pm$6.69 81.12$\pm$7.05 85.46
Table 10: Detailed results on CIFAR10-LT with ResNet50.
KNN@1 KNN@10 1%S LP
Method Head Mid Tail All Head Mid Tail All Head Mid Tail All
SimCLR 55.13 30.00 10.71 37.10 58.51 29.70 8.71 38.00 51.79 36.77 30.29 42.16
SDCLR 55.74 27.67 10.15 36.06 60.68 26.00 5.23 37.36 48.31 39.17 36.46 42.38
TS 57.23 30.26 13.14 38.36 60.41 29.53 10.14 38.86 60.41 40.38 33.57 45.18
TASE-800 59.26 31.84 10.92 39.54 64.11 30.24 8.31 40.26 54.16 40.78 37.85 45.48
TASE-1700 61.53 32.90 12.00 41.06 65.63 32.69 8.46 42.06 56.16 43.43 38.77 47.66
LT LP Full LP
Method Head        Mid        Tail         All Head        Mid         Tail         All
SimCLR 67.59       36.47       9.43       44.82 69.54       63.71       59.69       65.46
SDCLR 71.74       36.86       8.00       46.40 70.10       65.04       60.92       66.48
TS 70.67       38.85      10.29      47.26 72.53       56.27       63.69       67.82
TASE-800 71.16       39.88       9.54       47.82 72.74       66.41       62.46       68.30
TASE-1700 71.47       40.41       8.92       48.12 73.79       66.53       65.08       69.10
Table 11: Detailed results on ImageNet100-LT with ResNet50.TASE-800 denotes the result of training our method for 800 epochs of fair comparison. TASE-1700 train our method for 1700 epochs to verify scaling performance marked in italics.
Dataset Source Link
CIFAR10 https://www.cs.toronto.edu/ kriz/cifar-10-python.tar.gz
CIFAR100 https://www.cs.toronto.edu/ kriz/cifar-100-python.tar.gz
ImageNet http://image-net.org/download
Table 12: The source of the dataset we use.